Block A and B are moving toward each other. A has a mass of 2.6 kg and a velocity of 5.3m/s, while B has a mass of 6.7kg and a velocity of −7.6m/s. They suffer a completely inelastic collision. During the collision, how much kinetic energy is lost?


1
Expert's answer
2021-12-02T10:05:39-0500
"V(6.7+2.6)=5.3(2.6)-6.7(7.6)\\V=-4\frac{m}{s}\\\Delta K=0.5(5.3^2(2.6)+6.7(7.6)^2-(6.7+2.6)4^2)\\=156\ J"